OpenStudy (anonymous):

STATISTICS: Assume that the following data was collected for a chemical reaction where reactants A and B are reacting to form products C and D. As the products are formed, you measure their masses at two minute increments. Time (min) Amount of product(g) 2 3 6 5 7 7 8 10 10 13 12 17 14 21 16 26 18 34 20 50

OpenStudy (anonymous):

a) What is the equation for the model of best fit? Would it be log y=-0.076x*log x-o.1148? Or yhat=8.846-2.42x? Those are what I got. b) What does your model predict would have been the amount present at 5 minutes? I have no idea how to do this one.. c) At what time would 25.1 grams remain, according to your model? 15? d) Are the products formed at a constant rate over time? Explain. How would I determine this?

OpenStudy (anonymous):

Have you made a scatter plot yet?

OpenStudy (anonymous):

I can tell you this much for (a) a line with a negative slope is not going to fit the data at all.

OpenStudy (anonymous):

Yes, I have.

OpenStudy (anonymous):

I'm sure you can tell from the shape of it that it's an exponential relation. Fit an exponential model to it and then sub in x=5 to answer part (b). Put that point on your scatter plot to see that it fits in with the rest.

OpenStudy (anonymous):

How do I input that on my calculator (Ti-84)? Sorry, I feel so stupid >.>

OpenStudy (anonymous):

Your guess of 15min. for part (c) makes sense. After you have your equation, you can probably tighten up the precision of that.

OpenStudy (anonymous):

Eh, this stuff takes practice; don't sweat it. On the TI-84 input your data to L1 and L2 (STAT-EDIT) Then go to STAT-CALC and choose ExpReg L1,L2,Y1

OpenStudy (anonymous):

(you can choose Y1 by going to VARS-YVARS-Function-y1)

OpenStudy (anonymous):

I'm trying..

OpenStudy (anonymous):

I don't think I got it right.. It says 2.428*1.164 r2=.980 r=.990

OpenStudy (anonymous):

That's what I got too. You're just missing the exponent, x on the 1.164.

OpenStudy (anonymous):

\[y=2.4285(1.1646)^x\]

OpenStudy (anonymous):

for x=5, I got 5.2026

OpenStudy (anonymous):

I got 5.2036, but there might be round-off error somewhere.

OpenStudy (anonymous):

so that would be the answer to b?

OpenStudy (anonymous):

Yes, but make sure you include the unit of grams depending on how you are to submit your answers.

OpenStudy (anonymous):

thank you for explaining that to me :)